OpenStudy (anonymous):

Separate the real and imaginary parts from the following complex number: cos(a+ib) Also, what is the Magnitude and phase of the above complex number?

OpenStudy (phi):

I think you can use cos(x+y)= cos(x) cos(y) - sin(x) sin(y) and cos(ix)= cosh(x) sin(ix)= i sinh(x)

OpenStudy (anonymous):

Thats interesting! I should have supplied the hint now that I think about it. The hint that came with this problem is cos(X) = 1/2[e^(iX)+e^-(iX)]. I think I have all the answers besides phase, so if I could just get help for phase I would appreciate that!

OpenStudy (phi):

I think I have all the answers in that case, what did you get for the real and imag parts ?

OpenStudy (anonymous):

For real I got: [(0.5e^-b)*cos(a)+(0.5e^b)*cos(a)] For imaginary: [(0.5e^-b)*isin(a)-(0.5e^b)*isin(a)]

OpenStudy (phi):

I think you use for x+iy mag = sqrt( x^2 + y^2) phase = atan( y/x) It looks ugly for your expressions.

OpenStudy (anonymous):

Do you agree that I have the correct real/imaginary parts? I know, it looks real ugly huh lol

OpenStudy (phi):

yes, they look good. you could use the definition of cosh(b)= (e^b + e^-b)/2 and sinh(a) = (e^a - e^-a)/2 to write them as cos(a)cosh(b) - i sin(a) sinh(b) if you know about cosh and sinh

OpenStudy (anonymous):

Yes I think I can do that. Thank you very much phi!